Through (-2,-4) parallel to y=2x-3
Answer by Hawksfan(61)


y = 2x + 3 is of the form

y = mx + b where m = slope and b is the y intercept

when we want a line that is parallel, we keep the same slope and find a new
y intercept b
y= 2x + b (-2,-4)
we know x and y, solve for b
-4 = 2(-2) + b
-4 = -4 + b
b = 0
the new equation parallel to y=2x + 3 is y=2x +0 or y= 2x